OpenStudy (anonymous):

solve for x ? 8/x=x/11

OpenStudy (abb0t):

cross multiply to get \[x^2 = 88\]can you solve for x now?

OpenStudy (anonymous):

no

OpenStudy (abb0t):

square both sides to just get x You should get \[x = \pm \sqrt{a}\]

OpenStudy (abb0t):

a represents 88, obviously.

OpenStudy (jennychan12):

\[\frac{ 8 }{ x } = \frac{ x }{ 11 }\] cross multiply x^2 = 88 square root both sides \[x = \pm \sqrt{88} = \pm 2\sqrt{22} \approx 9.3808\]

OpenStudy (anonymous):

Jennychan12 has the correct answer

OpenStudy (jennychan12):

actually the others had the right work, they just didn't put the answer.